# DiMoDif: Discourse Modality-information Differentiation for Audio-visual Deepfake Detection and Localization

### v1

@@ 5.1 Temporal Forgery Localization
FakeAVCeleb [34] LAV-DF [10] AV-Deepfake1M [8] DFDC [20] KoDF [36]
ACC AP AUC ACC AP AUC ACC AP AUC ACC AP AUC ACC AP AUC
Training dataset FakeAVCeleb [34] 98.1 99.9 99.9 39.3 59.5 61.9 25.3 55.1 55.1 51.4 62.7 60.8 65.5 77.4 32.7
LAV-DF [10] 43.6 70.7 71.3 95.6 99.6 99.6 32.1 69.7 66.2 51.8 52.9 54.5 43.8 88.8 66.5
AV-Deepfake1M [8] 56.3 82.0 85.4 59.2 83.2 81.5 87.7 97.3 97.3 51.3 50.5 49.7 77.0 77.7 33.4
Table 7: Generalization performance on Deepfake detection task.
@@
5.2 Deepfake Detection
Tables 5, 5 and 6, present comparative results in terms of video-level Deepfake Detection (DFD) on LAV-DF [10], AV-Deepfake1M [8], and FakeAVCeleb [34], respectively. On LAV-DF, DiMoDif performs on-par with the state-of-the-art method UMMAFormer [76], achieving 99.6 AUC, better than the remaining competitive methods. On AV-Deepfake1M dataset, DiMoDif outperforms all state-of-the-art competitive methods by a significant absolute performance increase of +30.5% in terms of AUC [^7] Accuracy (ACC) is also reported by other papers; however, Codabench does not provide it for the test samples, thus we report the validation accuracy, which, even though is not comparable, still provides a sense on the level of uncalibrated threshold performance increase.. On FakeAVCeleb, DiMoDif outperforms all competitive methods on average achieving +4.1% absolute AP increase, and +2.8% absolute AUC increase, as well as in most forgery types. Finally, challenging real samples with normalized Levenshtein distance over 0.8 (cf. Section 1), are scored by DiMoDif with an average of 0.19, and 0.16 on LAV-DF, and AV-Deepfake1M, respectively, attesting to DiMoDif’s advantage wrt a naive distance-thresholding approach.
5.3 Generalization
This paper is the first to conduct generalization experiments on the used datasets. Table 7 presents DiMoDif’s cross-dataset performance on DFD. Training on FakeAVCeleb [34] leads to poor performance on LAV-DF, AV-Deepfake1M, and DFDC but to moderate performance on KoDF with 77.4 AP. Training on LAV-DF [10] leads to moderate performance on FakeAVCeleb and AV-Deepfake1M, low performance on DFDC but good performance on KoDF with 88.8 AP. Training on AV-Deepfake1M [8] dataset leads to high scores on FakeAVCeleb and LAV-DF with 85.4 AUC and 81.5 AUC respectively, moderate performance on KoDF with 77.7 AP, but low performance on DFDC. In general, low generalization is observed on DFDC, which can be partly explained by the discussion in Section 2.1. Our models focus on cross-modal information differences, while many DFDC videos do not contain talking or the talking person is different from the depicted one; thus, a visual-only face manipulation detector would likely lead to better results. Appendix C provides detailed generalization evaluation on FakeAVCeleb with scores per forgery type. Table 8 presents cross-dataset performance on TFL. Training on AV-Deepfake1M leads to acceptable performance on LAV-DF, in contrast to training on LAV-DF and evaluating on AV-Deepfake1M.
5.4 Ablation Analysis
(a) Window size $q$ (LAV-DF)
(b) Window size $q$ (AV-Deepfake1M)
(c) Feature pyramid use (LAV-DF)
(d) Feature pyramid use (AV-Deepfake1M)
(e) Transformer $\mathcal{T}$ size $(d,r,d\cdot u)$ (LAV-DF)
(f) Transformer $\mathcal{T}$ size $(d,r,d\cdot u)$ (AV-Deepfake1M)
Figure 3: Ablation and hyperparameter tuning analysis.
Figure 3 presents an ablation analysis that supports our methodological choices. For AV-Deepfake1M we report validation performance as conducting the ablations through Codabench would be impractical. These results pertain to TFL, while further ablations as well as the same analysis wrt DFD can be found in Appendix D. Figures 3(a) and 3(b) indicate that a small window size $q=15$ for Transformer’s $\mathcal{T}$ local attention is optimal, especially in contrast to attending on all tokens determined by $q=0$. Also, the feature pyramid scheme provides a small performance increase as indicated by Figures 3(c) and 3(d). The size of $\mathcal{T}$ plays crucial role in achieving maximum performance as shown in Figures 3(e) and 3(f), both in terms of number of layers $l$ and layer size determined by input dimension $d$, number of attention heads $r$, and internal dimension $d\cdot u$. Specifically, the larger the model the higher its performance.
6 Conclusions
In this work, we propose an audio-visual deepfake detection and localization framework that leverages cross-modal differences in machine perception of speech. It hinges on the assumption that the visual and audio signals of videos with real discourse coincide wrt information, in contrast to deepfakes that exhibit cross-modal incongruities. It considers a video and audio speech recognition feature extraction stage, and a Transformer encoder-based learning stage along with a combination of three loss functions that optimize frame-level detection, predicted interval overlap, and boundaries. An extensive evaluation study indicates the effectiveness of our framework reaching state-of-the-art performance and in several cases surpassing it. Our model is limited by the languages recognized by the used speech-recognition backbones, although it exhibits good generalization to Korean. Also, it assumes one continually visible speaking person, at least moderate video quality and face size to detect face landmarks, and requires both modalities for manipulation detection. Future work includes joint training on Deepfake Detection and Temporal Forgery Localization tasks to assess potential benefits and extension to multiple speakers.
7 Acknowledgments
Appendix A Objective function details
$\mathfrak{L}_{DFD}=-\frac{1}{2l}\sum_{m\in\{v,a\}}\sum_{j=1}^{l}y[m]\text{log}\hat{y}[j,m]+\\ (1-y[m])\cdot\text{log}(1-\hat{y}[j,m])$ (7)
where $j$ is layer index, $m$ is modality index, and $x[\cdot]$ is used to denote vector or matrix element access operation (array indexing).
  1. 1.
    $\mathfrak{L}_{focal}=\sum_{m\in\{v,a\}}\sum_{j=1}^{f}-\alpha_{t}^{m,j}(1-p_{t}^{m,j})^{\gamma}\text{log}(p_{t}^{m,j})$ (8)
    in which:
    $\alpha_{t}^{m,j}=\begin{cases}\alpha&\text{if }y^{m}_{j}=1\\ 1-\alpha&\text{otherwise}\end{cases}$ (9)
    $p_{t}^{m,j}=\begin{cases}\hat{y}^{m}_{j}&\text{if }y^{m}_{j}=1\\ 1-\hat{y}^{m}_{j}&\text{otherwise}\end{cases}$ (10)
    while $\alpha$ and $\gamma$ are hyperparameters.
  2. 2.
    $\mathfrak{L}_{DIoU}=\sum_{m\in\{v,a\}}\sum_{j=1}^{f}\Bigg{(}1.0-\text{IoU}(\hat{\mathbf{b}}^{m}_{j},\mathbf{b}^{m}_{j})+\\ \frac{\rho^{2}(\hat{b}^{m}_{j},b^{m}_{j})}{\kappa^{2}}\Bigg{)}\cdot\mathds{1}_{m,j}$ (11)
    where $j$ is frame index, $m$ is modality index, IoU denotes intersection over union, $\mathbf{b}^{m}_{j}=[s^{m}_{j},e^{m}_{j}]$ denotes a time interval, $b^{m}_{j}=0.5\cdot(e^{m}_{j}+s^{m}_{j})$ denotes the interval’s center, $\rho$ denotes the Euclidean distance, $\kappa$ denotes the length of the smallest enclosing interval covering both predicted and ground truth intervals, and $\mathds{1}_{m,j}$ is an indicator function that denotes if frame $j$ of modality $m$ is fake.
  3. 3.
    $\mathfrak{L}_{1}^{smooth}=\sum_{m\in\{v,a\}}\sum_{j=1}^{f}\frac{1}{2}\bigg{(}h(s^{m}_{j}-\hat{s}^{m}_{j})\\ +h(e^{m}_{j}-\hat{e}^{m}_{j})\bigg{)}\cdot\mathds{1}_{m,j}$ (12)
    in which $\mathds{1}_{m,j}$ is an indicator function that denotes if frame $j$ of modality $m$ is fake, and $h$ is defined as:
We combine the three loss functions with addition divided by the total number of fake visual and audio frames $p=\sum_{m\in\{v,a\}}\sum_{j=1}^{f}\mathds{1}_{m,j}$:
$\mathfrak{L}_{TFL}=\big{(}\mathfrak{L}_{focal}+\mathfrak{L}_{DIoU}+\mathfrak{L}_{1}^{smooth}\big{)}/p$ (14)
Appendix B Dataset class and split sizes
Table 9 provides details on dataset class and split sizes. Although, FakeAVCeleb paper [34] reports 10,000 FVFA and 9,000 FVRA samples, the download from the corresponding homepage https://sites.google.com/view/fakeavcelebdash-lab contains 10,835 and 9,709 respectively, resulting in 21.5K total number of samples instead of the commonly reported 20,000 size for this dataset. Documentation therein aslo claims “Since we apply the manual screening process on synthesized videos, the final video count is more than 20,000.”. Test metadata of AV-Deepfake1M have not been released but evaluation is enabled through Codabench.
FakeAVCeleb LAV-DF AV-Deepfake1M
train val test total train val test total train val test total
FVFA 8,678 1,052 1,105 10,835 19,090 7,701 6,369 33,160 186,344 14,515 - -
FVRA 7,760 975 974 9,709 19,271 7,820 6,452 33,543 186,597 14,304 - -
RVFA 399 49 52 500 19,088 7,709 6,373 33,170 186,573 14,286 - -
RVRA 17,766 1,535 1,199 20,500 21,254 8,271 6,906 36,431 186,666 14,235 - -
total 34,603 3,611 3,330 41,544 78,703 31,501 26,100 136,304 746,180 57,340 343,240 1,146,760
Appendix C Generalization on FakeAVCeleb
Table 10 provides a more detailed generalization evaluation on the FakeAVCeleb dataset with scores per forgery type, indicating the effectiveness of training on AV-Deepfake1M especially in some cases such as FVFA-WL and FVFA-FS reaching 95.6 AUC and 94.5 AUC, respectively. Training on LAV-DF also results in good generalization especially in FVFA-FS case with 90.0 AUC.

### v2

@@ 5.6 Generalization
FakeAVCeleb LAV-DF AVD1M*
AP AUC AP AUC AP AUC
Training dataset FakeAVCeleb 99.99 99.71 93.10 84.47 77.91 54.00
LAV-DF 99.69 90.25 99.94 99.84 88.46 70.40
AV-Deepfake1M 99.69 90.69 94.98 86.30 99.72 99.18
Table 9: Generalization on DFD. * Validation set.
5.7 Real-world Analysis
We also present an analysis of DiMoDif’s performance in the wild. Initially, we compiled a collection of 204 Internet videos, across several languages (incl. Out-Of-Distribution $\mathtt{OOD}$) and varying lengths ($\mathtt{max}$ 46, $\mathtt{median}$ 1), with 75 of them being real and 129 of them being fake. Fakes were selected for being popular deepfakes, supporting conspiracy theories, or spreading political misinformation, and real videos for containing discourse (e.g., news anchoring). Next, we developed a pre-processing pipeline to address model’s limitations. Chunking is considered for accepting any-length videos, while frame groups with small faces (e.g. depicting crowds or wide shots), no faces, or small duration (¡2s) are discarded. We utilize DiMoDif trained on AV-Deepfake1M (TFL), and evaluate the ratio of fake to valid video parts, $\mathtt{c}=\mathtt{fake}/\mathtt{valid}$, as the model occasionally flags small parts of authentic videos as deepfakes. DiMoDif achieves 82.1 AP, indicating robust detection of real-world cases, with linear processing complexity $\mathcal{O}(\mathtt{[sec]})$. Figure 7 illustrates the distribution of $\mathtt{c}$ for real vs. fake and decision examples. For reference, RealForensics [29] achieves 76.4 AP on the same set. Additionally, although the developed pipeline is tailored to real-world cases it excels on evaluation benchmarks, i.e., achieves 99.1 AP on FakeAVCeleb, 93.5 AP on LAV-DF*, and 90.6 AP on AV-Deepfake1M* (*1K random unseen samples). Further details are provided in supplementary material.
Figure 7: Distribution of coverage $\mathtt{c}$ for in-the-wild real vs. fake and examples of decisions. Red parts are flagged as fake.
6 Conclusions
In this work, we propose an audio-visual deepfake detection and localization framework that leverages cross-modal differences in machine perception of speech. It hinges on the assumption that the visual and audio signals of videos with real discourse coincide wrt information, in contrast to deepfakes that exhibit cross-modal incongruities. It considers a visual and audio speech recognition feature extraction stage, and a hierarchical cross-modal fusion learning stage along with a combination of three loss functions that optimize frame-level detection, predicted interval overlap, and boundaries. An extensive evaluation study indicates the effectiveness of our framework exhibiting state-of-the-art performance.
Limitations: DiMoDif’s language coverage is constrained by its speech recognition backbones although it was found to generalize well to Korean (cf. supplementary material) and in-the-wild $\mathtt{OOD}$ examples, while performance may be degraded by modality misalignment in real-world samples. Additionally, the trained model assumes a single, consistently visible speaker, sufficient video quality for facial landmark detection, and the availability of both visual and audio modalities. Development for the real-world analysis has addressed in part some of these limitations through video pre-processing. We aim to further improve in-the-wild applicability and extend functionality to multi-speaker scenarios.
7 Acknowledgments
Appendix A Objective function details
$\mathfrak{L}_{DFD}=\mathfrak{L}_{ce}=-\frac{1}{2l}\sum_{m\in\{v,a\}}\sum_{\lambda=1}^{l}y[m]\text{log}\hat{y}[\lambda,m]+\\ (1-y[m])\cdot\text{log}(1-\hat{y}[\lambda,m])$ (7)
where $\lambda$ is layer index, $m$ is modality index, and $x[\cdot]$ is used to denote vector or matrix element access operation (array indexing).
  1. 1.
    $\mathfrak{L}_{f}=\sum_{m\in\{v,a\}}\sum_{\phi=1}^{f}-\alpha_{t}^{m,\phi}(1-p_{t}^{m,\phi})^{\gamma}\text{log}(p_{t}^{m,\phi})$ (8)
    in which:
    $\alpha_{t}^{m,\phi}=\begin{cases}\alpha&\text{if }y^{m}_{\phi}=1\\ 1-\alpha&\text{otherwise}\end{cases}$ (9)
    $p_{t}^{m,\phi}=\begin{cases}\hat{y}^{m}_{\phi}&\text{if }y^{m}_{\phi}=1\\ 1-\hat{y}^{m}_{\phi}&\text{otherwise}\end{cases}$ (10)
    while $\alpha$ and $\gamma$ are hyperparameters.
  2. 2.
    $\mathfrak{L}_{d}=\sum_{m\in\{v,a\}}\sum_{\phi=1}^{f}\Bigg{(}1.0-\text{IoU}(\hat{\mathbf{b}}^{m}_{\phi},\mathbf{b}^{m}_{\phi})+\\ \frac{\rho^{2}(\hat{b}^{m}_{\phi},b^{m}_{\phi})}{\kappa^{2}}\Bigg{)}\cdot\mathds{1}_{m,\phi}$ (11)
    where $\phi$ is frame index, $m$ is modality index, IoU denotes intersection over union, $\mathbf{b}^{m}_{\phi}=[s^{m}_{\phi},e^{m}_{\phi}]$ denotes a time interval, $b^{m}_{\phi}=0.5\cdot(e^{m}_{\phi}+s^{m}_{\phi})$ denotes the interval’s center, $\rho$ denotes the Euclidean distance, $\kappa$ denotes the length of the smallest enclosing interval covering both predicted and ground truth intervals, and $\mathds{1}_{m,\phi}$ is an indicator function that denotes if frame $\phi$ of modality $m$ is fake.
  3. 3.
    $\mathfrak{L}_{1}^{s}=\sum_{m\in\{v,a\}}\sum_{\phi=1}^{f}\frac{1}{2}\bigg{(}h(s^{m}_{\phi}-\hat{s}^{m}_{\phi})+\\ h(e^{m}_{\phi}-\hat{e}^{m}_{\phi})\bigg{)}\cdot\mathds{1}_{m,\phi}$ (12)
    in which $\mathds{1}_{m,\phi}$ is an indicator function that denotes if frame $\phi$ of modality $m$ is fake, and $h$ is defined as:
We combine the three loss functions with addition divided by the total number of fake visual and audio frames $p=\sum_{m\in\{v,a\}}\sum_{\phi=1}^{f}\mathds{1}_{m,\phi}$:
$\mathfrak{L}_{TFL}=\big{(}\mathfrak{L}_{f}+\mathfrak{L}_{d}+\mathfrak{L}_{1}^{s}\big{)}/p$ (14)
Appendix B Dataset class and split sizes
Table 10 provides details on dataset class and split sizes. Although, FakeAVCeleb paper [36] reports 10,000 FVFA and 9,000 FVRA samples, the download from the corresponding homepage https://sites.google.com/view/fakeavcelebdash-lab contains 10,835 and 9,709 respectively, resulting in 21.5K total number of samples (incl. 500 RVFA & 500 RVRA) instead of the commonly reported 20,000 size for this dataset. Documentation therein aslo claims “Since we apply the manual screening process on synthesized videos, the final video count is more than 20,000.”. Test metadata of AV-Deepfake1M have not been released but evaluation is enabled through Codabench.
Appendix C Further Ablations
Figure 8 presents the ablations on LAV-DF (TFL) corresponding to the ablations presented in the main manuscript on AV-Deepfake1M, namely for window size $q$, feature pyramid use, and Transformer $\mathcal{T}$ size. The conclusions drawn from these ablations on both datasets agree. Figure 9 then presents further TFL task ablations on both LAV-DF and AV-Deepfake1M. Specifically, Figures 9(a) and 9(b) indicate little to no sensitivity to the focal loss’ hyperparameter $\alpha$. Figures 9(c) and 9(d) indicate that the use of a learning rate scheduler is beneficial however the type of scheduler is not significantly affecting the results. Next, Figure 10 presents ablations on FakeAVCeleb, LAV-DF, and AV-Deepfake1M wrt the DFD task. The results indicate that on DFD a window size of $q=15$ is optimal, but performance is not increased with feature pyramids nor with different learning rate schedules. Also, medium-sized models achieve maximum performance.
Appendix D Interpretability
Figures 11 and 12 present further interpretability plots for 8 real and 8 fake video samples from LAV-DF. Cosine similarity at the frame-level reveals DiMoDif’s decision reasoning in these cases. However, there are some rare cases in which DiMoDif correctly identified the modified parts although cross-modal similarity is high, such as in the example presented in Figure 13. In such cases DiMoDif is forced to successfully identify alternative manipulation features sacrificing interpretability.
Appendix E Robustness
Robustness to visual distortions: Figure 14 illustrates the performance of DiMoDif under unseen visual perturbations (our model has been trained without any visual augmentations) commonly present in real-world scenarios. We follow the evaluation protocol of [54, 23, 30] which consider the FakeAVCeleb test set and directly source the implementations of visual perturbations from https://github.com/EndlessSora/DeeperForensics-1.0. Specifically, videos are subject to saturation, contrast, block-wise, Gaussian noise, Gaussian blur, JPEG compression, and video compression, while we also compute the average robustness across perturbations. DiMoDif clearly outperforms RealForensics [29] under all perturbation scenarios except video compression at intensity levels 4 and 5. Also, our method outperforms AVFF [54] under saturation, contrast, block-wise, Gaussian blur, and JPEG compression at all levels of intensity, and is outperformed by AVFF under Gaussian noise and video compression at all levels of intensity. Wrt average AUC DiMoDif slightly outperforms AVFF and is much more robust than RealForensics. Notably, DiMoDif exhibits exceptional robustness wrt AP on average achieving 99.7 at intensity level 5, indicating near-perfect detection of positive samples despite significant visual distortions. In approximately 0.06% of cases, high-intensity Gaussian noise (level 5) prevented landmark detection, making lip-reading impossible and thus precluding the application of our method.
Robustness to audio distortions: Figure 15 illustrates the performance of DiMoDif under unseen audio perturbations (our model has been trained without any audio augmentations) commonly present in real-world scenarios. We follow the evaluation protocol of [54] that considers the FakeAVCeleb test set and implement the audio perturbations using torchaudio, pysndfx, and pydub Python libraries as indicated therein. Specifically, the videos are subject to Gaussian noise, pitch shift, reverberence, and audio compression at 5 levels of intensity. DiMoDif achieves unprecedented levels of robustness to audio perturbations with 99.3 average AUC across all perturbations at intensity level 5. This fact indicates that audio speech recognition is not affected by such perturbations and the corresponding features are effective under any perturbation level scenario. In addition, DiMoDif outperforms AVFF under all perturbation scenarios and at all intensity levels.
[…]
Appendix F Generalization
We evaluate DiMoDif (trained on FakeAVCeleb, LAV-DF, and AV-Deepfake1M) on the test set of DFDC [21] and a randomly selected 6K-sized sample from KoDF [38] proportionally spanning all synthesis methods and real samples.
Table 12 presents the corresponding results. Training on FakeAV-Celeb [36] leads to poor performance on DFDC but to high performance on KoDF with 98.8 AP. Training on LAV-DF [10] leads to almost random performance on DFDC but high performance on KoDF with 97.9 AP. Training on AV-Deepfake1M [8] dataset leads to random performance on DFDC but high performance on KoDF with 98.9 AP. In general, very high generalization performance is observed on KoDF, although Korean are not recognized by the (Ma 2022) backbones, indicating the learning of cross-language features by DiMoDif. Also, low generalization is observed on DFDC, which can be partly explained by the discussion in Section 2.2. Our models focus on cross-modal speech-related information differences, while many DFDC videos contain multiple talking persons [^12] E.g., DFDC/test/aalyqplqns.mp4, moving persons [^13] E.g., DFDC/test/chyfcacfjr.mp4 or wide shots [^14] E.g., DFDC/test/ljnjhvpnhq.mp4" rendering lip-reading hard, do not contain talking [^15] E.g., DFDC/test/aarpyivfys.mp4 or the talking person is different from the depicted one [^16] E.g., DFDC/test/mfjljhqhca.mp4; thus, visual-only face manipulation detectors lead to better generalization in this dataset.
Table 13 presents cross-dataset performance on TFL wrt all AP@{0.5, 0.75, 0.90, 0.95} and AR@{100, 50, 30, 20, 10, 5} metrics. Cross-dataset performance on TFL is shown to be much more challenging than on DFD. Also, the results are even worse when training on LAV-DF in contrast to AV-Deepfake1M which contains half-sized fake parts, thus presenting a more challenging objective, while also being much larger.
Table 14 provides a more detailed generalization evaluation on the FakeAVCeleb dataset with scores per forgery type, indicating the effectiveness of training on AV-Deepfake1M reaching 99.2 AP & 95.5 AUC (AVG-FV) and 93.9 AP & 95.5 AUC (RVFA). Training on LAV-DF also results in good generalization reaching 98.5 AP & 91.2 AUC (AVG-FV) and 81.3 AP & 76.8 AUC (RVFA). Note, that this is not a cross-manipulation analysis as the models trained on LAV-DF and AV-Deepfake1M datasets have seen all manipulation types (RVRA, RVFA, FVFA, FVRA). Thus, metrics of the 2nd and 3rd rows are comparable but the 1st row (showing cross-manipulation performance) is only provided for reference and is not comparable with the other two.
Appendix G Real-world analysis
LABEL:tab:real_world_info provides further information on the set of videos used for the real-world analysis. Specifically, we provide the public URL (if exists), the spoken language, the duration of the video, the processing (by DiMoDif) time, and the Coverage score $\mathtt{c}$. Language has been automatically retrieved through https://speechbrain.github.io/. Finally, 16 of the videos do not have public URLs.
Appendix H Error analysis
Figure 16 presents two examples of DiMoDif’s misclassifications. Specifically, in the real video shown in fig. 16(a), a non-verbal vocalization (panting imitation) between 6.3s and 7.3s is incorrectly interpreted as inconsistency, leading DiMoDif to assign a high fake probability to those frames. In contrast, fig. 16(b) features a visual-only manipulated segment from 0.3s to 0.78s; here, despite a slight reduction in cross-modal inconsistency, the high fidelity of the utterance deceives the model.
FakeAVCeleb LAV-DF AV-Deepfake1M
train val test total train val test total train val test total
FVFA 6,996 561 3,278 10,835 19,090 7,701 6,369 33,160 186,344 14,515 - -
FVRA 6,363 470 2,876 9,709 19,271 7,820 6,452 33,543 186,597 14,304 - -
RVFA 326 24 150 500 19,088 7,709 6,373 33,170 186,573 14,286 - -
RVRA 13,689 22 160 13,871 21,254 8,271 6,906 36,431 186,666 14,235 - -
total 27,374 1,077 6,464 34,915 78,703 31,501 26,100 136,304 746,180 57,340 343,240 1,146,760
(a) LAV-DF: Window size $q$
(b) LAV-DF: Feature Pyramid (FP) use
(c) LAV-DF: $\mathcal{T}$’s size; $l$ $(d,r,d\cdot u)$
Figure 8: Ablation and hyperparameter tuning analysis on LAV-DF (TFL).
(a) LAV-DF: Focal loss hyperparmeter $\alpha$
(b) AV-Deepfake1M: Focal loss hyperparmeter $\alpha$
(c) LAV-DF: Learning rate scheduler
(d) AV-Deepfake1M: Learning rate scheduler
Figure 9: Ablation and hyperparameter tuning analysis on LAV-DF and AV-Deepfake1M (TFL).
(a) FakeAVCeleb: Window size $q$
(b) LAV-DF: Window size $q$
(c) AV-Deepfake1M: Window size $q$
(d) FakeAVCeleb: Feature Pyramid (FP) use
(e) LAV-DF: Feature Pyramid (FP) use
(f) AV-Deepfake1M: Feature Pyramid (FP) use
(g) FakeAVCeleb: $\mathcal{T}$’s size; $l$ $(d,r,d\cdot u)$
(h) LAV-DF: $\mathcal{T}$’s size; $l$ $(d,r,d\cdot u)$
(i) AV-Deepfake1M: $\mathcal{T}$’s size; $l$ $(d,r,d\cdot u)$
(j) FakeAVCeleb: Learning rate scheduler
(k) LAV-DF: Learning rate scheduler
(l) AV-Deepfake1M: Learning rate scheduler
Figure 10: Ablation and hyperparameter tuning analysis on FakeAVCeleb, LAV-DF, and AV-Deepfake1M (DFD).
(a) Real #1: LAV-DF/test/000000.mp4
(b) Real #2: LAV-DF/test/000001.mp4
(c) Real #3: LAV-DF/test/000005.mp4
(d) Real #4: LAV-DF/test/000012.mp4
(e) Real #5: LAV-DF/test/000016.mp4
(f) Real #6: LAV-DF/test/000032.mp4
(g) Real #7: LAV-DF/test/000036.mp4
(h) Real #8: LAV-DF/test/000040.mp4
Figure 11: Cross-modal cosine similarity on DiMoDif’s representations of 8 real examples.
(a) Fake #1: LAV-DF/test/000003.mp4
(b) Fake #2: LAV-DF/test/000008.mp4
(c) Fake #3: LAV-DF/test/000011.mp4
(d) Fake #4: LAV-DF/test/000014.mp4
(e) Fake #5: LAV-DF/test/000015.mp4
(f) Fake #6: LAV-DF/test/000018.mp4
(g) Fake #7: LAV-DF/test/000019.mp4
(h) Fake #8: LAV-DF/test/000030.mp4
Figure 12: Cross-modal cosine similarity on DiMoDif’s representations of 8 fake examples.
Figure 13: Non-interpretable fake example (LAV-DF/test/000017.mp4).
Figure 14: Robustness to unseen visual perturbations. AVFF and RealForensics performance taken from [54]’s supplementary material.
Figure 15: Robustness to unseen audio perturbations. AVFF performance taken from [54]’s supplementary material.
Table 11: Text prediction robustness of DiMoDif’s feature extraction backbones (Ma 2022) [46], to visual and audio distortions for the real video sample: FakeAVCeleb/RealVideo-RealAudio/Caucasian (European)/men/id00055/00120.mp4. First row presents predictions on the original video while the next ones use versions of it with visual or audio distortions as input. $\mathcal{V}$: transcript from visual input, $\mathcal{A}$: transcript from audio input.
Dist. Lev. Transcript
- - $\mathcal{V}$: you know i played with growing up and always had great ability great talent from a young age
$\mathcal{A}$: know i played with growing up and he always had great ability great talent from a young age
Visual distortions
Saturation 1 $\mathcal{V}$: you know i played with growing up and always had great ability great talent from a young age
2 $\mathcal{V}$: how i played with growing up and always had great ability great talent from a young age
3 $\mathcal{V}$: you know i played with growing up and always had great ability great talent from a young age
4 $\mathcal{V}$: yeah i played with growing up and always had great ability great talent from a young age
5 $\mathcal{V}$: yeah i played with growing up and always had great ability great talent from a young age
Contrast 1 $\mathcal{V}$: yeah i played with growing up and always had a great ability great talent from a young age
2 $\mathcal{V}$: how i played with growing up and always had great ability great talent from a young age
3 $\mathcal{V}$: played with growing up and always had great ability great talent from a young age
4 $\mathcal{V}$: you know i played with growing up and always had great ability great talent from a young age
5 $\mathcal{V}$: yeah i played with harold growing up and he always had great ability great talent from a young age
Block-wise 1 $\mathcal{V}$: yeah i played with growing up and always had great ability great talent from a young age
2 $\mathcal{V}$: yeah i played with growing up and always had great ability great talent from a young age
3 $\mathcal{V}$: yeah i played with growing up and always had great ability great talent from a young age
4 $\mathcal{V}$: yeah i played with growing up and always had great ability great talent from a young age
5 $\mathcal{V}$: yeah i played with growing up and always had great ability great talent from a young age
Gauss. noise 1 $\mathcal{V}$: i played with carol growing up and he always had a great ability great talent from a young age
2 $\mathcal{V}$: i played with her growing up and it was a great ability great talent from a young age
3 $\mathcal{V}$: everybody was scared of growing up with this great ability great talent from a young age
4 $\mathcal{V}$: our plan was to start growing up and you can see how i beat the grass down from the high
5 $\mathcal{V}$: -
Gauss. blur 1 $\mathcal{V}$: i played with growing up and always had great ability great talent from a young age
2 $\mathcal{V}$: how i played with carol growing up and he always had a great ability great talent from a young age
3 $\mathcal{V}$: how i played with carol growing up and he always has a great ability great talent from a young age
4 $\mathcal{V}$: how i played with carol growing up he always has great ability great talent from a young age
5 $\mathcal{V}$: now i played with carol growing up he always had great ability great talent from a young age
JPEG compr. 1 $\mathcal{V}$: you know i played with growing up and always had great ability great talent from a young age
2 $\mathcal{V}$: i played with growing up and always had great ability great talent from a young age
3 $\mathcal{V}$: now i’m here with the care of growing up and i always like to see the great ability to raise hands for without hands
4 $\mathcal{V}$: i will play with you the character i’ll be happy once i turn to you i’ll play you as i grew up in high
5 $\mathcal{V}$: there were pictures of his car flipped up so he said to the red bricks at the right angle that’s
Video compr. 1 $\mathcal{V}$: yeah i played with growing up and always had great ability great talent from a young age
2 $\mathcal{V}$: how i played with how i’ve grown up and always had the right ability the right talent from a young age
3 $\mathcal{V}$: a little bit we’ve now grown up and we’re starting to create building a great town from a young age
4 $\mathcal{V}$: now i’ll be able to fire from the camera you know adjust it to the front but the right hand will be high
5 $\mathcal{V}$: i’ll make sure it’s just a narrow volume count it will take you to the right place and the right hand will be high
Audio distortions
Gauss. noise 1 $\mathcal{A}$: know i played with growing up and he always had great ability great talent from a young age
2 $\mathcal{A}$: know i played with growing up and he always had great ability great talent from a young age
3 $\mathcal{A}$: know i played with growing up and he always had great ability great talent from a young age
4 $\mathcal{A}$: know i played with growing up and he always had great ability great talent from a young age
5 $\mathcal{A}$: know i played with growing up and he always had great ability great talent from a young age
Pitch shift 1 $\mathcal{A}$: know i played with growing up and he always had great ability great talent from a young age
2 $\mathcal{A}$: yeah i played with growing up and he always had great ability great talent from a young age
3 $\mathcal{A}$: you know papers cars growing up and it was a great time for the young age
4 $\mathcal{A}$: yeah played of gaff growing up and he was a great and it’s a great kind of for a young age
5 $\mathcal{A}$: -
Reverberence 1 $\mathcal{A}$: know i played with growing up and he always had great ability great talent from a young age
2 $\mathcal{A}$: know i played with growing up and he always had great ability great talent from a young age
3 $\mathcal{A}$: know i played with growing up and he always had great ability great talent from a young age
4 $\mathcal{A}$: know i played with growing up and he always had great ability great talent from a young age
5 $\mathcal{A}$: know i played with growing up and he always had great ability great talent from a young age
Audio compr. 1 $\mathcal{A}$: know i played with growing up and he always had great ability great talent from a young age
2 $\mathcal{A}$: know i played with growing up and he always had great ability great talent from a young age
3 $\mathcal{A}$: know i played with growing up and he always had great ability great talent from a young age
4 $\mathcal{A}$: know i played with growing up and he always had great ability great talent from a young age
5 $\mathcal{A}$: know i played with growing up and he always had great ability great talent from a young age
Test dataset
DFDC [21] KoDF [38]
AP AUC AP AUC
Training dataset FakeAVCeleb [36] 60.8 58.9 98.8 69.9
LAV-DF [10] 52.9 54.5 97.9 56.7
AV-Deepfake1M [8] 50.5 49.8 98.9 72.8
Table 12: Generalization on Deepfake detection task.